# Anomaly Detection and RFI Classification with Unsupervised Learning in Narrowband Radio Technosignature Searches

### v1

@@ 4.1 Random Forest Analysis
The benefit of this approach comes from demonstrating in a general sense that no one feature is significantly more important than the others. If one were, it could cause the HDBSCAN algorithm to cluster on the basis of it alone. If a feature does turn out to be unusually important, measures can be taken to reduce its importance, such as downscaling its range in pre-processing.
@@ 4.3 Principal Component Analysis
PCA has been used as a pre-processing step in other clustering frameworks when a lower-dimensional (Sooknunan et al., 2021; Biswas et al., 2023) and/or decorrelated (Matchev et al., 2022) feature space is needed. However, the features found by PCA do not share equal importance. Instead, they contain different fractions of the total variance of the original data set, permitting the reconstruction of its broadest structures with only the few highest-variance features (“principal components”), but requiring more for finer details.
@@
The curve in Figure 7 is fairly smooth and does not jump suddenly from a low variance fraction to a fraction near 1.0. More than 99% of the variance can be retained with as few as 10 components, but in a parameter space where much RFI looks very alike, even 1% can be significant for clustering and anomaly detection. For lack of any obvious cutoff point, we keep all 13 original features and forgo PCA as a pre-processing step for our analysis. Using the full 13 principal components may also be a valid approach for reasons of de-correlation, since the principal components are linearly independent by definition, but we do not find any correlations in our feature space strong enough to merit PCA for their mitigation.
@@ 5 Results
We also tested our method’s recovery rate of injected synthetic signals. We simulated 100 drifting narrowband signals with drift rates between $-3$ and $3$ Hz s-1 and SNRs between 1,000 and 10,000, and injected them into the “on” pointings of cadence observations of M 31, a total of 300 hit injections. This is a different set of synthetic signal injections than the 100 seed signals used to identify clusters of interest in each batch. Of the 300 hit injections, 296 were identified by the FindDoppler hit search, corresponding to a possible 98 out of the 100 injected events. The GLOBULAR clustering process erroneously removed 32 injected hits, leaving 264.
@@ 6 Conclusions
In this work, we demonstrate an application of the HDBSCAN clustering algorithm to technosignature searches, where it significantly reduces the false positive rate while maintaining an injected signal recovery rate only slightly lower than current search methods. HDBSCAN does not scale easily to data sets of the volume regularly seen in radio astronomy, but we mitigate the deleterious effects of upscaling with a novel iterative batching and pooling scheme. This implementation significantly reduces both noise and compute time over an approach that runs HDBSCAN once on all data points.
@@
Active learning presents another potentially useful tool for refinement. Through a front-end interface, a user can manually relabel an active learning model’s first guess at signal classifications, one signal at a time, to improve future classifications. By preferentially selecting signals that occur at the boundaries of clusters, the user need only relabel a small fraction of the data set for their feedback to dramatically improve the model’s classification overall. This method has already been applied to great effect by Lochner & Bassett (2021) in their package Astronomaly, which was originally designed for anomaly detection in galaxy images but is extendable to other survey types and data products, including in the radio, as demonstrated by Andersson et al. (2024). As Etsebeth et al. (2024) show, as few as $\sim 10^{5}$ user labels are necessary for effective classification and anomaly detection in a data set of 3,884,404 galaxy images, easily achievable by one person in a few hours. Similar principles are likely to apply for dynamic spectra in comparably sized radio surveys.
@@ Acknowledgements
The Breakthrough Prize Foundation funds the Breakthrough Initiatives, which manages Breakthrough Listen. The Green Bank Observatory facility is supported by the National Science Foundation, and is operated by Associated Universities, Inc., under a cooperative agreement. The authors thank Verlon Etsebeth and Peter Ma for helpful discussions, and BJB is grateful to the 2024 Berkeley SETI REU participants for their support.

### v2

@@ 4 Application to Small Data Sets
If the analysis is performed on only a single batch’s worth of signals, the computational cost of iterating many runs to optimize these hyperparameters is low. However, in the case that significantly broader ranges of values need to be tested, or that the batch size is significantly larger than we use here, utilities such as CORE-SG (Neto et al., 2022) and MustaCHE (Neto et al., 2018) may be useful for efficiently iterating over $n_{\text{pts}}$ values and for visualizing the output, respectively.
@@ 4.1 Random Forest Analysis
The benefit of this approach comes from demonstrating in a general sense that no one feature is significantly more important than the others. If one were, it could cause the HDBSCAN algorithm to cluster on the basis of it alone. If a feature does turn out to be unusually important, measures can be taken to reduce its importance, such as downscaling its range in preprocessing.
@@ 4.3 Principal Component Analysis
PCA has been used as a preprocessing step in other clustering frameworks when a lower-dimensional (Sooknunan et al., 2021; Biswas et al., 2023) and/or decorrelated (Matchev et al., 2022) feature space is needed. However, the features found by PCA do not share equal importance. Instead, they contain different fractions of the total variance of the original data set, permitting the reconstruction of its broadest structures with only the few highest-variance features (“principal components”), but requiring more for finer details.
@@
The curve in Figure 7 is fairly smooth and does not jump suddenly from a low variance fraction to a fraction near 1.0. More than 99% of the variance can be retained with as few as 10 components, but in a parameter space where much RFI looks very alike, even 1% can be significant for clustering and anomaly detection. For lack of any obvious cutoff point, we keep all 13 original features and forgo PCA as a preprocessing step for our analysis. Using the full 13 principal components may also be a valid approach for reasons of decorrelation, since the principal components are linearly independent by definition, but we do not find any correlations in our feature space strong enough to merit PCA for their mitigation.
@@ 5 Results
We also tested our method’s recovery rate of injected synthetic signals. We simulated 100 drifting narrowband signals with drift rates between $-3$ and $3$ Hz s-1 and SNRs between 1,000 and 10,000, and injected them into the “on” pointings of cadence observations of M 31, a total of 300 hit injections. This is a different set of synthetic signal injections than the 100 seed signals used to identify clusters of interest in each batch. Of the 300 hit injections, 296 were identified by the FindDoppler hit search, corresponding to a possible 98 out of the 100 injected events. From the identified injections, the GLOBULAR clustering process erroneously removed 32 hits, leaving 264.
@@ 6 Conclusions
In this work, we demonstrate an application of the HDBSCAN clustering algorithm to technosignature searches, where it significantly reduces the false positive rate while maintaining an injected signal recovery rate only slightly lower than current search methods. HDBSCAN does not scale easily to data sets of the volume regularly seen in radio astronomy, but we mitigate the deleterious effects of upscaling with an iterative batching and pooling scheme. This implementation significantly reduces both noise and compute time over an approach that runs HDBSCAN once on all data points.
@@
Active learning presents another potentially useful tool for refinement. Through a front-end interface, a user can manually relabel an active learning model’s first guess at signal classifications, one signal at a time, to improve future classifications. By preferentially selecting signals that occur at the boundaries of clusters, the user need only relabel a small fraction of the data set for their feedback to dramatically improve the model’s classification overall. This method has already been applied to great effect by Lochner & Bassett (2021) in their package Astronomaly, which was originally designed for anomaly detection in galaxy images but is extendable to other survey types and data products, including in the radio, as demonstrated by Andersson et al. (2024). As Etsebeth et al. (2024) show, as few as $\sim 10^{5}$ user labels are necessary for effective classification and anomaly detection in a data set of 3,884,404 galaxy images, easily achievable by one person in a few hours. Similar principles are likely to apply for dynamic spectra in comparably sized radio surveys, and will be investigated in detail in future work.
@@ Acknowledgements
The Breakthrough Prize Foundation funds the Breakthrough Initiatives, which manages Breakthrough Listen. The Green Bank Observatory facility is supported by the National Science Foundation, and is operated by Associated Universities, Inc., under a cooperative agreement. CP was funded as a participant in the Berkeley SETI Research Center Research Experience for Undergraduates Site, supported by the National Science Foundation under Grant No. 2244242. The authors thank the anonymous referee for valuable comments and Verlon Etsebeth and Peter Ma for helpful discussions, and BJB is grateful to the 2024 Berkeley SETI REU participants for their support.